Bob's father is 3 times old as Bob. 4 years ago, he was 4 times older. How old is Bob?
Answer by fractalier(2101)

Always assign variables in age problems in the NOW.
Let B be Bob's age NOW.
Let F be the father's age NOW.
From the problem,
F = 3B
F - 4 = 4(B - 4)
Substituting, we get
3B - 4 = 4B - 16
-B = -12
and
B = 12